Solve each by substitution 3x -4y =-22 and -3x +y = o
Answer by KnightOwlTutor(293)

-3x+y=0
y=3x
In the second equation replace 3x with y and solve for y
3x -4y =-22
y-4y=-22
-3y=-22
divide both sides by -3
y=22/3

y=3x
22/3=3x
divide both sides by 3
22/9=x
Let's double check by plugging into the equation
-3(22/9)+22/3=0

3(22/9)-88/3=-22